OpenStudy (anonymous):

g(x)=2x^2-3x+10; g'(x)=4x-3 How would you find g'(x) when g(x) =15?

OpenStudy (anonymous):

didn't you already find g'(x)?

OpenStudy (anonymous):

Or if you mean g'(15), You would find the derivative of \(g(x)\) and evaluate this at 15. Since you have most genourously already obtained the derivative, just evaluate it haha: \[\eqalign{ &g'(x)=4x-3 \\ &g'(15)=4(15)-3 \\ &g'(15)=60-3\\ &g'(15)=57 \\ }\]

OpenStudy (anonymous):

This was how the question was stated. It seemed weird to me.

OpenStudy (ddcamp):

Solve g(x) = 15 for x 2x² - 3x + 10 = 15 → 0 = 2x² - 3x - 5 → x= 2.5, -1 Then, plug those values of x into g'(x)

OpenStudy (campbell_st):

substitute g(x) with 15 and you get \[15 = 2x^2 - 3x + 10\] solve for x when you get x, subsitute into g'(x)

OpenStudy (anonymous):

Haha that was my third guess

OpenStudy (anonymous):

Well is g(x)=15, the derivative is really easy to calculate. The derivative of a constant is 0. So g'(x)=0.